OpenStudy (anonymous):

find an equation of the line through the pair points. write the eqaution in the form Ax+By=C (-6,3) and (4,0)

myininaya (myininaya):

y=mx+b we can find m \[m=\frac{3-0}{-6-4}=\frac{3}{-10}=-\frac{3}{10}\] we can find b we know m and we know a x and a y (x,y)=(4,0) \[0=-\frac{3}{10}*4+b\] \[0=\frac{-12}{10}+b\] \[b=\frac{12}{10}=\frac{6}{5}\] \[y=\frac{-3}{10}*x+\frac{6}{5}\] now lets write it in the form Ax+By=C \[10y=-3x+12\] \[3x+10y=12\]

OpenStudy (anonymous):

m = (y1-y2) / (x1-x2) = (3-0)/(-6-4) = 3/-10 m= -3/10 y = -3/10x + c 3 = -3/10(-6) + c ...substitute in and solve for C 3 = 18/10 + c c = 3 - 18/10 c = (30-18)/10 c = 12/10 y = -3/10x + 12/10 multiply all by 10 10y = -3x + 12 therefore 3x + 10y = 12